OpenStudy (anonymous):

If x^2+y^2=25, what is the value of d^2y/dx^2 at the point (4,3)?

OpenStudy (anonymous):

Do you need the answer or explanation?

OpenStudy (anonymous):

d?

OpenStudy (anonymous):

It's for derivatives right?

OpenStudy (anonymous):

This may go wrong somewhere, but I think this is how it goes: x^2+y^2=25 y^2=-x^2+25 y=sqrt(-x^2+25)

OpenStudy (anonymous):

That's where we start. Now, we need to find the first derivative. We do this by using the formula for the derivatives of composite functions. \[dy/dx=dy/du.du/dx\]

OpenStudy (anonymous):

So, y=sqrt(u) u=(-x^2+25) y'=1/2(u^-1/2) (y' means derivative of y) u'=-2x Therefore, dy/dx=1/2(u^-1/2)(-2x) =-x(-x^2+25)^-1/2

OpenStudy (anonymous):

\[d ^{2}y / dx ^{2}=-(-x ^{2}+25)^{-1/2}-x(-1/2)(-x ^{2}+25)^{-3/2}\]

OpenStudy (dumbcow):

i get \[d^{2}y/dx = \frac{2x^{2} -25}{\sqrt{25-x^{2}}(25-x^{2})}\] plug in x=4

OpenStudy (anonymous):

Now, substituting 4 in, we should get the answer. -1/3+2/27 =-7/27

OpenStudy (dumbcow):

is that the same as yours?? i'm off by a neg

OpenStudy (anonymous):

uhhh, just the negative at the front of my equation is different